A F-15 fighter plane is view from the ground; the angle enclosed by the length of plane is theta= 0.01744 radians at the point of observation. The Technical specifications of F-15 indicate that the length of the plane is 19.45 m. What is the height (r) of the plane from the ground?
1
Expert's answer
2020-05-12T10:23:54-0400


In the figure we see that


"r=\frac{l}{2\text{tan}(\theta/2)}=\frac{19.45}{2\text{tan}[(0.01744 \cdot180/\pi)/2]}=1115\text{ m}."